OpenStudy (anonymous):

Find the radius of a cylinder with a volume of 118.79 feet cubed and a height of 5 feet. A)3.45 B)1.94 C)2.52 D)2.75

OpenStudy (lalaly):

\[\text{volume of cylinder}=\pi r^2 h\]\[118.79=\pi r^2 \times 5 \]now solve for r

Parth (parthkohli):

^ AWESOME MEMBER ON THE SITE ^_^

OpenStudy (lalaly):

Lol .. nothing compared to you xD

OpenStudy (anonymous):

@lalaly : how do i solve for r ?

Parth (parthkohli):

Isolate r.

Parth (parthkohli):

xP

OpenStudy (lalaly):

divide by 5pi ... then take the squareroot of both sides ...

Parth (parthkohli):

lol you won

OpenStudy (lalaly):

haha

OpenStudy (anonymous):

dont understand /:

OpenStudy (lalaly):

do u know how to make r subject of the formula

OpenStudy (lalaly):

hmm ... u are asked to find the radius .. which is r ... so u have to make r on one side all alone to find its value... ill walk u through it... from the equation i gave u for the volume what do u think u should do to make r alone

OpenStudy (anonymous):

3.14(5) ?

OpenStudy (lalaly):

yeah so divide both sides by that to get rid of them ... so u will have\[\frac{118.79}{5 \times 3.14}=r^2\]now find that value

OpenStudy (anonymous):

i got 7.56624

OpenStudy (lalaly):

yeah thats right,,, now thats the value of r^2 ... what do u think we sould do to make it r?

OpenStudy (anonymous):

ummm divide that by 5 ?

OpenStudy (lalaly):

noo where did the 5 come from You have\[\large{r^2=7.56624}\]how can u make it r= a number

OpenStudy (anonymous):

idk i told you i was bad at math. but do you divide that by pi?

OpenStudy (lalaly):

u take the square root of both sides

OpenStudy (anonymous):

2.75 is my answer (:

OpenStudy (lalaly):

yep thats right

OpenStudy (anonymous):

thank you so much. if my teachers went step by step like that i wouldnt of been in this situationlol

OpenStudy (lalaly):

lol its ok just keep trying and ull learn ... Ur welcome and goodluck:)